The operation @ is defined as @(a) = a^2 + 4a -4, for any real number a. Find the value(s) of a for which @(a)= a.
Answer by stanbon(48545)

The operation @ is defined as @(a) = a^2 + 4a -4, for any real number a. Find the value(s) of a for which @(a)= a.
-------------
If @(a)=a then,
a^2+4a-4=a
Subtract "a" from both sides to get:
a^2+3a-4=0
Factor to get:
(a+4)(a-1)=0
Solve for "a" to get:
a=-4 or a=1
=============
Cheers,
Stan H.